40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Feb 25 15:30:29 2013, yuta, Update, Alignment, Y arm locked, both colors 
    Reply  Mon Feb 25 17:58:28 2013, rana, Update, Alignment, Y arm locked, both colors 
       Reply  Mon Feb 25 22:42:32 2013, yuta, Summary, Alignment, current IFO situation IPANGdrift.png
    Reply  Mon Feb 25 21:25:14 2013, yuta, Update, Alignment, both arms locked in IR QUAD1_1045890588.png
       Reply  Tue Feb 26 01:49:27 2013, Jenne, Update, Locking, PRMI locked 
          Reply  Tue Feb 26 02:04:51 2013, Koji, Update, Locking, PRMI locked 
             Reply  Tue Feb 26 11:55:11 2013, Manasa, Update, Locking, FPMI locked 
                Reply  Tue Feb 26 15:32:47 2013, Jenne, Update, Locking, FPMI locked 
             Reply  Wed Feb 27 01:09:53 2013, yuta, Update, Locking, PRMI sideband/carrier locked 6x
                Reply  Wed Feb 27 02:44:39 2013, Jenne, Update, Locking, DRFPMI flashes 
                   Reply  Wed Feb 27 20:10:43 2013, rana, Update, Locking, DRFPMI flashes 
Message ID: 8164     Entry time: Mon Feb 25 22:42:32 2013     In reply to: 8158
Author: yuta 
Type: Summary 
Category: Alignment 
Subject: current IFO situation 

[Jenne,Yuta]

Both arms are aligned starting from Y green.
We have all beams unclipped except for IPANG. I think we should ignore IPANG and go on to PRMI locking and FPMI locking using ALS.
IPANG/IPPOS and oplev steering mirrors are kept un-touched after pumping until now.

Current alignment situation:
 - Yarm aligned to green (Y green transmission ~240 uW)
 - TT1/TT2 aligned to Yarm (TRY ~0.86)
 - BS and Xarm alined to each other (TRX ~ with MI fringe in AS)
 - X green is not aligned yet
 - PRMI aligned

Current output beam situation:
 IPPOS - Coming out clear but off in yaw. Not on QPD.
 IPANG - Coming out but too high in pitch and clipped half of the beam. Not on QPD.
 TRY   - On PD/camera.
 POY   - On PD.
 TRX   - On PD/camera.
 POX   - On PD.
 REFL  - Coming out clear, on camera (centered without touching steering mirrors).
 AS    - Coming out clear, on camera (centered without touching steering mirrors).
 POP   - Coming out clear, on camera (upper left on camera).


Oplev values:

Optic    Pre-pump(pit/yaw)    PRFPMI aligned(pit/yaw)
ITMX    -0.26 /  0.60         0.25 /  0.95
ITMY    -0.12 /  0.08         0.50 /  0.39
ETMX    -0.03 / -0.02        -0.47 /  0.19
ETMY     0.37 / -0.62        -0.08 /  0.80
BS      -0.01 / -0.18        -1    /  1 (almost off)
PRM     -0.34 /  0.03        -1    /  1 (almost off)


 All values +/- ~0.01. So, oplevs are not useful for alignment reference.

OSEM values:
Optic    Pre-pump(pit/yaw)    PRFPMI aligned(pit/yaw)
ITMX    -1660 / -1680        -1650 / -1680
ITMY    -1110 /   490        -1070 /   440
ETMX     -330 / -5380         -380 / -5420
ETMY    -1890 /   490        -1850 /   430
BS        370 /   840          360 /   800
PRM      -220 /  -110         -310 /  -110


  All values +/- ~10.
  We checked that if there's ~1200 difference, we still see flash in Watec TR camera. So, OSEM values are quite good reference for optic alignment.

IPANG drift:
  On Saturday, when Rana, Manasa, and I are trying to get Y arm flash, we noticed IPANG was drifting quite a lot in pitch. No calibration is done yet, but it went off the IPANG QPD within ~1 hour (attached).
  When I was aligning Yarm and Xarm at the same time, TRY drifted within ~1 hour. I had to tweak TT1/TT2 mainly in yaw to keep TRY. I also had to keep Yarm alignment to Y green. I'm not sure what is drifting so much. Suspects are TT2, PR2/PR3, Y arm and Y green.

  I made a simple script(/opt/rtcds/caltech/c1/scripts/Alignment/ipkeeper) for keeping input pointing by centering the beam on IPPOS/IPANG using TT1/TT2. I used this for keeping input pointing while scanning Y arm alignment to search for Y arm flash this weekend (/opt/rtcds/caltech/c1/scripts/Alignment/scanArmAlignment.py). But now we have clipped IPANG.


So, what's useful for alignment after pumping?:

  Optic alignment can be close by restoring OSEM values. For input pointing, IPPOS/IPANG are not so useful. Centering the beam on REFL/AS (POP) camera is a good start. But green works better.

Attachment 1: IPANGdrift.png  49 kB  | Hide | Hide all
IPANGdrift.png
ELOG V3.1.3-